OpenStudy (anonymous):

Find the distance from (2,-1) to the line y=2x+3. The answer is 8/5*root5 by the way, but I want to know how to solve. I used distance formula. sqrt ( (x-2)^2 + (y+1)^2) sqrt ( (x-2)^2 + (2x+3+1)^2) sqrt ( (x-2)^2 + (2x+4)^2) sqrt (5x^2 + 12x + 20) When I enter this on my calculator to solve for x, it cannot do it. What should I do next?

OpenStudy (anonymous):

you don't use X

OpenStudy (precal):

What kind of calculator are you using? Are you using the solve feature?

OpenStudy (anonymous):

Yes I am usin the solve feature.

OpenStudy (precal):

Why don't you use graph paper instead?

OpenStudy (anonymous):

I have to use algebraic methods and show my work.

OpenStudy (precal):

graphing on graph paper is a way to show your work

OpenStudy (anonymous):

Is there a way to answer the problem algebraically?

OpenStudy (anonymous):

Well, the line to the other linen willl be perpendicular to it

OpenStudy (anonymous):

Assuming you want the shortest distance

OpenStudy (precal):

using graph paper is a valid method, then you can count the number of units

OpenStudy (anonymous):

Find the equation of a line perpendicular to y=2x+3 and use the distance formula on the intersection

OpenStudy (precal):

You could use a point from the equation. I would graph it to decide what point to use. You need two ponts for the distance formula

OpenStudy (mertsj):

Have you studied the formula for finding the perpendicular to a line through a point?

OpenStudy (anonymous):

Well, I'm in Calculus right now, but I don't know how to continue this problem.

OpenStudy (precal):

Yep calculus would study that Do the normal line

OpenStudy (anonymous):

Wouldn't the normal slope be -1/2 ?

OpenStudy (anonymous):

-1/2*

OpenStudy (anonymous):

\[y+1=\frac{-1}{2}(x-2)\]

OpenStudy (anonymous):

\[y=-\frac{1}{2}x\]

OpenStudy (mertsj):

You could put the equation in the form Ax+By=C=0 and use this formula: \[d=\frac{|Am+Bn+C|}{\sqrt{A ^{2}+B ^{2}}}\] The point is (m,n)

OpenStudy (anonymous):

Now set 2x+3=-(1/2)x to find the intersection, then use the distance formula

OpenStudy (anonymous):

x = -6/5

OpenStudy (anonymous):

Get y now, and use this point in your distance formula

OpenStudy (precal):

Mertsj what formula did you post? Do you know the name of it? Just curious

OpenStudy (mertsj):

Perpendicular distance from a Point to a Line

OpenStudy (precal):

ok I have not seen that one before. Good to know

OpenStudy (anonymous):

Okay I now have two methods of approaching this problem. Thank you all for helping me. :)